Find the volume of the solid obtained by rotating the region bounded by y = x^2 +1 and y = 9 – x^2 about y = -1.

1
Expert's answer
2022-03-21T09:16:12-0400

Let’s go to new coordinates:

"\~x=x";

"\~y=y+1".

Now the volume of the solid can be obtained by rotating the region bounded by "\~y=\~x^2+2" and "\~y=10-\~x^2" about "\~y=0".

Interval of integration:

"\~x^2+2= 10-\~x^2"

"\~x=\pm2"

"V=\int_{-2}^2 \pi((10-\~x^2)^2-(\~x^2+2)^2)d\~x="

"\int_{-2}^2 \pi(96-24\~x^2)d\~x=\frac12 \int_{0}^2 \pi(96-24\~x^2)d\~x="

"\frac12\pi (96\~x-8\~x^3)|_0^2=\pi(96-8\cdot4)=64\pi"

Answer: "V=64\pi" .